OpenStudy (anonymous):

Please Help! Giving Medal! A psychology student conducts an ESP experiment in which the subject attempts to identify the number that turns up when an unseen die is rolled. The die is rolled six times. On the basis of chance alone, find the probability that the subject identifies the number correctly four of the six times

OpenStudy (anonymous):

exactly four out of six?

OpenStudy (anonymous):

\[\binom{6}{4}(\frac{1}{6})^4(\frac{5}{6})^2\] and a calculator

OpenStudy (anonymous):

btw \(\binom{6}{4}=\frac{6\times 5}{2}=3\times 5=15\)

OpenStudy (anonymous):

Alright, Thank You =) I Really appreciate it